Find the value of "y" so that the line passing through the two points has the given slope: (5,7) , (8,y) m= 4/3
Answer by princess08(3)

m=(y-7)/(8-5)
4/3=(y-7)/3
(4/3)*3=y-7
4=y-7
11=y
y=11.